### v2

# These authors contributed equally

Abstract
Counterfactuals have been established as a popular explainability technique which leverages a set of minimal edits to alter the prediction of a classifier. When considering conceptual counterfactuals on images, the edits requested should correspond to salient concepts present in the input data. At the same time, conceptual distances are defined by knowledge graphs, ensuring the optimality of conceptual edits. In this work, we extend previous endeavors on graph edits as counterfactual explanations by conducting a comparative study which encompasses both supervised and unsupervised Graph Neural Network (GNN) approaches. To this end, we pose the following significant research question: should we represent input data as graphs, which is the optimal GNN approach in terms of performance and time efficiency to generate minimal and meaningful counterfactual explanations for black-box image classifiers?
Keywords:
Counterfactual Explanations Black-box Explanations Scene Graphs Graph Neural Networks Graph Autoencoders.
1 Introduction
In the era of large and complex neural models, ensuring trust between them and humans becomes a critical issue. Explainability literature suggests a variety of methods to probe neural models behaviors, either requiring access to their inner workings [11, 29] or not [6, 7, 10]. There is also ample discussion regarding the nature of the features involved in a human-understandable explanation; for example, low-level features e.g. pixel-related characteristics (brightness, contrast) may be unable to provide a meaningful explanation to the end user, despite being informative for a neural model [27]. This observation is applicable in the counterfactual explanation scenario, suggesting that semantics are fundamentally essential to meaningful counterfactual explanations [4]. In the meanwhile, both high-level semantics as well as low-level features can be expressed in the same mathematical format (e.g. numerical vectors), allowing the transition to semantically rich explanation systems [6]. In the interest of explaining visual classification systems, and with respect to recent related work [2, 6, 7, 10, 11, 29], we base our current study on conceptual counterfactual explanations.
Figure 1: Left - person in front of parked car (safe). Right - person in front of moving cars on the highway (unsafe). The relationship between the person, the car and the highway is critical for the transition to the counterfactual class.
With regard to the model accessibility an explainability system can have, most related counterfactual explanation methods belong to the white-box category [2, 11, 29], offering insights tailored to the model under investigation. However, in the emergence of powerful, yet proprietary models, such as ChatGPT [23] and GPT-4 [24] which are currently accessible only through APIs, black-box explanations resurface as more viable and universal solutions, able to approach a model by scrutinizing their outputs. Under these constrains, we stay within the black-box setting of conceptual counterfactuals [6, 7, 10].
We also acknowledge the importance of intra-concept roles, as associations between the same concepts via roles may result in different interpretations of situations (Figure 1). Thus, we advocate image representation via scene graphs. To this end, we extend the framework of [7] to accommodate a larger spectrum of Graph Machine Learning algorithms to serve conceptual counterfactual explanations on scene images. We retain the brute-force calculation of Graph Edit Distance (GED)[28] as the ground truth counterfactual measure, under which the closest scene graph in terms of GED belonging to a different class is considered as the ”gold” counterfactual graph. In order to overcome the computational complexity of GED calculation, we employ Graph Machine Learning algorithms, such as Graph Kernels, Graph Autoencoders (GAEs) and supervised Graph Neural Networks (GNNs), which are proven to be lightweight and significantly faster than the deterministic GED calculation, even when optimizations [14] are employed. Graph edits serving as counterfactual explanations are instructed via concept distances in existing hierarchies, in our case being WordNet [21]. The outline of our approach is presented in Figure 2. Overall, we explore the capabilities of Graph Machine Learning algorithms, contributing to the following:
  • We prove quantitatively and qualitatively that both unsupervised (GAEs) and supervised (GNNs) methods can provide meaningful and accurate conceptual counterfactuals, viewing models under explanation as black-boxes.
  • We demonstrate the trade-offs between unsupervised vs supervised approaches towards counterfactual explanations, and provide insights on the advantages of each method.
2 Related work
Counterfactual Explanations have been established as an effective explanation method [31] should certain requirements, such as actionability and feasibility be met [26]. These principles are transferred in the field of counterfactual visual explanations, which aim to explain the behaviors of image classifiers. Most of these approaches leverage pixel-level edits, denoting regions that should be minimally changed [11, 13, 29] and often proceed to actually generate the targeted region by employing generative models [3, 5, 8, 33].
Since semantics are inherently tied to human-interpretable counterfactuals [4], a diverging line of work suggests moving towards conceptual rather than pixel-level counterfactuals. The notion of explainable concepts was presented in [2], where fault-lines are utilized to drive concept edits. [1] produce conceptual counterfactuals to meaningfully explain model misclassifications. However, these approaches require some degree of access to the internals of the classifier. Minimal concept edits in a black-box fashion were proposed in [10], paving the way for knowledge-based model-agnostic visual counterfactuals. Some limitations were resolved in subsequent works, suggesting an indirect [6] or direct [7] incorporation of roles into concepts, while underlining the importance of correctly selecting data and representations to obtain proper explanations. The same idea is modified to explain hallucinations of image generation models rather than classifiers [20], highlighting the significance of universal black-box explainers.
3 Method
The focus of our current study centers on determining the minimal semantic alteration necessary for an image $i$ to transition from being classified as A to B. Each image $i$ in our dataset $I$ is represented by a scene graph $G_{i}$, encapsulating semantic information, therefore forming an explanation dataset {$i,G_{i}$}, as previously outlined in [10]. Utilizing a reference image $i_{r}$ and candidate scene graphs $G_{i_{c}}$, the Graph Edit Distance (GED) is computed using an optimization approach to identify the most efficient path for semantic transformation between $G_{i_{r}}$ and $G_{i_{c}}$, guided by predefined graph edit operations (Replacement, Deletion, Insertion). Semantic distances, drawn from the well-crafted WordNet [21] hierarchy, guide the choice of edits, ensuring meaningful transitions while penalizing irrelevant replacements. To accelerate GED calculations, we employ optimization techniques similar to [7], utilizing the Volgenant-Jonker (VJ) algorithm [14]. Finally, to obtain the counterfactual candidate $G_{i_{cc}}$ of the original scene graph $G_{i_{r}}$, we repeat the optimized GED calculation for each candidate $G_{i_{c}}$, where $c\in C=I\setminus\{i_{r}\}$, ensuring that the selected closest $G_{i_{cc}}$ belongs to class $B\neq A$, while $G_{i_{r}}$ belongs to class $A$, formulated as:
$i_{cc}=\operatorname*{arg\,min}_{i_{c}\in C}\left(GED\left(G_{i_{r}},G_{i_{c}}\right)\right),\;\;\;\;\;G_{i_{r}}\in A,\;G_{i_{c}}\in B\neq A$ (1)
Figure 2: Outline of our evaluation framework: Given scene graphs of classes $A\neq B$, a graph model embeds them in a low-dimensional space, allowing the retrieval of the closest graphs $G_{A}$, $G_{B}$ from which we extract counterfactual edits.
3.1 The importance of Graph Machine Learning
Considering a dataset with $N$ graphs, calculating the counterfactual $G_{i_{cc}}$ for every $G_{i_{r}}$ requires $O(N^{2})$ GED operations, resulting in a prohibitive calculation time, despite using optimizations. In order to accelerate this process, we leverage Graph Machine Learning algorithms to circumvent the exhaustive calculation of $N^{2}$ GEDs, accepting an unavoidable performance sacrifice.
Graph Kernels is our first step towards testing graph matching [17]. Kernels in Machine Learning are responsible for mapping the given data in high-dimensional spaces, revealing complex patterns that cannot be observed in the original dimensionality of the input data, therefore allowing the application of linear algorithms to solve non-linear problems. Graph kernels quantify the degree of similarity between subgraphs in polynomial time by employing kernel functions, thus contributing to the overall structural similarity between two given graphs. In our experiments, we will compare different graph kernel approaches.
Graph Neural Networks (GNNs) in the supervised setting comprise architectures such as Graph Convolution Network (GCN) [15], Graph Attention Network (GAT) [30], and Graph Isomorphism Network (GIN) [32]. They are incorporated in a Siamese Network to learn similarity on input instances [19], demonstrating state-of-the-art results even when trained on a smaller subset of $N^{2}/j$ labeled graph pairs ($N$ is the number of graphs in the complete dataset, $j$ a positive integer). During training, these GNN variants learn the relationships between counterfactual graph pairs utilizing ground truth samples calculated from Eq. 1. Embedding representations on new graphs are extracted using the trained GNNs, translating graph similarity cues to distances in the embedding space, therefore allowing retrieval of counterfactual graphs $G_{i_{cc}}$ for each $G_{i_{r}}$.
Graph Autoencoders (GAEs) [16] arise as faster and more versatile solutions by eliminating the need for labeled samples, even though their metric performance often ranks lower than GNNs. GAEs implement an encoder-decoder structure, with the encoder being tasked to produce embedding representations; these representations are useful for retrieval tasks, since similarity in the embedding space corresponds to actual graph similarity. Thus, the need of GED calculations within the training process is completely avoided, dropping the computational complexity to $O(N)$ in comparison to the $O(N^{2})$ operations required in the supervised case ($N$ is the number of graphs). The encoder and decoder modules are constructed using any of the aforementioned GNN variants (same GNN modules for the encoder and the decoder when applicable).
In all cases, we retain the NP-hard GED scores for ground truth comparison to evaluate the performance of different models on the ground-truth rank.
4 Experiments
Table 1: Counterfactual retrieval results with Supervised and Unsupervised GNN architectures. Underlined cells indicate best results per graph model (kernel, supervised, unsupervised). Bold denotes best results overall per dataset split.
Model NDCG (binary) Precision (P) (binary) Precision (P)
@4 @2 @1 @4 @2 @1 @4 @2 @1
VG-DENSE
WL kernel 0.287 0.198 0.138 0.186 0.128 0.076 0.164 0.125 0.076
SP kernel 0.334 0.251 0.194 0.232 0.174 0.122 0.154 0.145 0.122
RW kernel 0.0 0.0 0.0 0.0 0.0 0.0 0.001 0.0 0.0
NH kernel 0.282 0.192 0.131 0.124 0.086 0.05 0.088 0.073 0.05
GS kernel 0.259 0.166 0.103 0.018 0.008 0.008 0.022 0.007 0.008
Superv. GAT 0.342 0.26 0.204 0.422 0.288 0.164 0.312 0.25 0.164
Superv. GIN 0.339 0.256 0.2 0.344 0.242 0.158 0.252 0.192 0.158
Superv. GCN 0.407 0.333 0.283 0.492 0.37 0.248 0.358 0.292 0.248
GAE GAT 0.281 0.191 0.13 0.108 0.078 0.048 0.075 0.075 0.048
GAE GIN 0.281 0.191 0.13 0.132 0.098 0.052 0.082 0.078 0.052
GAE GCN 0.287 0.197 0.137 0.102 0.078 0.052 0.071 0.06 0.052
VGAE GAT 0.298 0.21 0.15 0.196 0.138 0.082 0.139 0.12 0.082
VGAE GIN 0.304 0.216 0.157 0.202 0.146 0.09 0.141 0.119 0.09
VGAE GCN 0.287 0.198 0.137 0.174 0.118 0.066 0.122 0.101 0.066
GFA GAT 0.302 0.214 0.155 0.182 0.14 0.086 0.122 0.109 0.086
GFA GIN 0.309 0.223 0.164 0.21 0.148 0.096 0.14 0.122 0.096
GFA GCN 0.163 0.165 0.158 0.289 0.2 0.139 0.12 0.104 0.068
ARVGA GAT 0.3 0.213 0.153 0.188 0.14 0.084 0.132 0.114 0.084
ARVGA GIN 0.317 0.232 0.174 0.218 0.164 0.106 0.135 0.124 0.106
ARVGA GCN 0.295 0.207 0.146 0.184 0.138 0.076 0.122 0.106 0.076
VG-RANDOM
WL kernel 0.306 0.219 0.16 0.166 0.124 0.096 0.13 0.108 0.096
SP kernel 0.292 0.204 0.144 0.11 0.09 0.064 0.079 0.068 0.064
RW kernel 0.0 0.0 0.0 0.0 0.0 0.0 0.007 0.0 0.0
NH kernel 0.303 0.216 0.157 0.168 0.116 0.092 0.135 0.116 0.092
GS kernel 0.0 0.0 0.0 0.002 0.0 0.0 0.05 0.0 0.0
Superv. GAT 0.35 0.268 0.213 0.382 0.294 0.176 0.294 0.245 0.176
Superv. GIN 0.327 0.243 0.186 0.308 0.228 0.144 0.245 0.205 0.144
Superv. GCN 0.369 0.29 0.236 0.424 0.3 0.2 0.295 0.246 0.2
GAE GAT 0.294 0.205 0.146 0.132 0.104 0.07 0.092 0.085 0.07
GAE GIN 0.297 0.209 0.15 0.114 0.096 0.07 0.08 0.076 0.07
GAE GCN 0.291 0.203 0.143 0.114 0.1 0.066 0.079 0.078 0.066
VGAE GAT 0.305 0.218 0.159 0.16 0.122 0.09 0.103 0.1 0.09
VGAE GIN 0.304 0.217 0.158 0.152 0.118 0.088 0.11 0.101 0.088
VGAE GCN 0.303 0.216 0.157 0.156 0.122 0.088 0.101 0.089 0.088
GFA GAT 0.292 0.203 0.143 0.142 0.108 0.076 0.107 0.094 0.076
GFA GIN 0.298 0.211 0.151 0.142 0.11 0.082 0.11 0.093 0.082
GFA GCN 0.297 0.209 0.149 0.14 0.112 0.082 0.106 0.095 0.082
ARVGA GAT 0.305 0.218 0.159 0.156 0.13 0.092 0.111 0.103 0.092
ARVGA GIN 0.298 0.21 0.15 0.152 0.118 0.084 0.116 0.099 0.084
ARVGA GCN 0.306 0.22 0.161 0.154 0.13 0.094 0.113 0.097 0.094
Table 2: Average Number of node, edge and total edits, as well as Average Top-1 GED for counterfactual instances retrieved for VG-DENSE and VG-RANDOM.
Model VG-DENSE VG-RANDOM
Node $\downarrow$ Edge $\downarrow$ Total $\downarrow$ GED $\downarrow$ Node $\downarrow$ Edge $\downarrow$ Total $\downarrow$ GED $\downarrow$
WL kernel 5.24 11.668 16.908 130.146 13.44 11.946 25.386 192.824
SP kernel 4.964 12.078 17.042 129.334 14.144 12.386 26.53 208.096
RW kernel 9.162 20.006 29.168 217.684 17.814 22.36 40.174 385.726
NH kernel 5.348 11.264 16.612 141.18 12.568 11.714 24.282 190.376
GS kernel 6.444 16.748 23.192 178.434 17.216 15.344 32.56 272.744
Superv. GAT 5.278 10.892 16.17 108.636 12.81 11.978 24.788 159.324
Superv. GIN 5.106 10.766 15.872 116.174 12.792 11.394 24.186 173.72
Superv. GCN 4.942 10.312 15.254 105.194 12.392 11.84 24.232 159.356
GAE GAT 5.504 11.848 17.352 143.976 11.952 11.426 23.378 197.626
GAE GIN 5.4 12.002 17.402 143.182 12.554 11.738 24.292 201.638
GAE GCN 5.178 11.27 16.448 143.506 12.134 11.384 23.518 202.638
VGAE GAT 5.03 11.4 16.43 130.124 12.884 12.376 25.26 197.226
VGAE GIN 4.918 10.97 15.888 131.514 12.692 12.042 24.734 199.34
VGAE GCN 5.114 11.702 16.816 132.056 12.91 12.392 25.302 195.938
GFA GAT 5.07 11.564 16.634 131.582 12.988 12.412 25.4 198.78
GFA GIN 4.856 10.786 15.642 129.856 12.916 12.218 25.134 200.974
GFA GCN 5.03 11.432 16.462 132.046 12.696 12.338 25.034 201.038
ARVGA GAT 5.062 11.482 16.544 133.476 12.508 12.192 24.7 196.946
ARVGA GIN 4.894 10.912 15.806 128.944 12.234 11.794 24.028 198.654
ARVGA GCN 5.048 11.316 16.364 132.294 12.464 12.088 24.552 198.46
In this experimental section, we aim to provide a thorough comparison between different kernel/GAE/GNN architectures as components of the counterfactual retrieval process. We examine them both in terms of their adherence to ground truth rankings, as well as their ability to provide minimal edits regarding both quantity and cost. Throughout this analysis, we inherently contrast supervised and unsupervised GNNs, weighing the trade-offs between employing more efficient versus accurate models and quantifying the resultant incurred damages.
4.1 Experimental setup
Data
In all our experiments, we select Visual Genome (VG) [18] as the experimental dataset, due to its abundance in real-life scenes accompanied by manually annotated scene graphs. In addition, VG offers an effortless connection with WordNet, as synset annotations of objects and roles are also provided by annotators. Since graph edges are important in our analysis, we regard two density-based dataset splits for training [7]: the first one, called VG-DENSE, includes graphs with many interconnections and fewer isolated nodes. The second one called VG-RANDOM does not impose any restrictions on scene graph density. Both splits contain $N$=500 graphs. As VG does not contain ground truth scene labels we assign them using a pre-trained PLACES-365 classifier [34].
Graph Models
In our study, we employ various graph kernels [22] including Shortest Path (SP), Weisfeiler-Lehman (WL), Neighborhood Hash (NH), Random Walk (RW), and Graphlet Sampling (GS), implemented using the GraKel library. For supervised Graph Neural Network (GNN) models, we utilize single-layer siamese Graph Convolutional Network (GCN), Graph Attention Network (GAT), and Graph Isomorphism Network (GIN) as described in [7], trained over 50 epochs with a batch size of 32 and Adam optimizer. Regarding Graph Autoencoders (GAEs), we implement vanilla GAE, Variational GAE (VGAE) [16], and Adversarially Regularized Variational Graph Autoencoder (ARVGA) [25], alongside a variation of Graph Feature Autoencoder (GFA) [12]. GAE and VGAE are both based on the generic auto-encoder structure, but differ in terms of loss function. ARVGA employs basic concepts from Generative Adversarial Networks to further boost the regularization of the latent embeddings produced by VGAE. Finally, GFA utilizes the Feature Decoder (whose goal is to predict the feature matrix $X$), alongside the original VGAE Inner-Product Decoder (whose goal is to predict the adjacency matrix $A$). All these models incorporate single-layer GCN, GAT, and GIN architectures in the encoder (and the decoder for GFA), employing AdamW optimizer with default parameters and a learning rate of 0.001, trained for 20 epochs with a batch size of 32.
GNNs and GAEs were implemented using PyTorch Geometric [9]. In all cases, cosine similarity was used as a measure for embedding similarity. For each scene graph in the dataset acting as the query, we retrieve the top $k$ closest graphs with different PLACES-365 labels using a pretrained PLACES classifier with a ResNet50 backbone, resulting in slightly improved labels compared to [7]. We then evaluate whether the counterfactual graph, determined by the ground truth Graph Edit Distance (GED), is found within this ranked list of $k$ items.
Evaluation
Information retrieval metrics are used for evaluation, specifically Precision (P@k) and Normalized Discounted Cumulative Gain (NDCG@k). P@k returns the percentage of the relevant items, when considering the first $k$ items in the rank. We also employ the ”binary” variant from [7], where only the top-1 Graph Edit Distance (GED) item is relevant, in line with the counterfactual logic. Binary NDCG@k compares the rank of the top-1 counterfactual returned by an algorithm to the ideal ground truth rank obtained by GED. While NDCG@k lacks direct interpretability, it’s widely used for comparing retrieval system performance. Our evaluation focuses on $k$=1, 2, 4 ranks, once again emphasizing proximity to the top-1 result. Additional metrics assess the proposed edits’ minimality by GNNs, including average node, edge, and overall edits, alongside average GED of counterfactuals. Lower values indicate better model performance, with these metrics considered relative to each other.
4.2 Quantitative Results
In Table 1, we present ranking results regarding counterfactual retrieval using all proposed graph models: graph kernels, supervised GNNs and unsupervised GAEs. For both dataset subsets, all supervised models were trained using 70K scene graph pairs extracted from $N$=500 graphs, while all unsupervised models were trained on the same 500 graphs (without pairing them).
Kernels reveal distinct performance patterns, with some closely rivaling unsupervised autoencoders. Notably, in VG-DENSE, SP and WL kernels excel, while in VG-RANDOM, WL and NH show superior performance. However, SP and NH metrics falter in VG-RANDOM and VG-DENSE, respectively, underscoring the significance of edge density. Conversely, kernels like RW consistently fail to retrieve counterfactual graphs, indicating their limitations in capturing graph structure. Supervised GNN models, particularly GCN, outperform kernels in ranking metrics, benefiting from optimized learning on ground truth GED. Unsupervised autoencoders exhibit more variability, with ARVGA GIN excelling in VG-DENSE and VG-RANDOM splits, underlining the importance of adversarial regularization for competitive performance.
In Table 2, we present results on the average number of edits, a common metric for evaluating counterfactual explanation systems, alongside the average top-1 counterfactual GED. In VG-DENSE experiments, the superior performance of supervised GCN is evident in edit numbers. While total edits appear similar across models, GFA GIN emerges as the second-best performer, indicating the limitations of relying solely on edit counts without considering their cost (GED). Comparing supervised GCN and GFA GIN, a significant gap in GED underscores their semantic divergence. Results for VG-RANDOM exhibit more variability, with GAE GAT leading in total edits. Considering both edit counts and GED provides insights into explanation minimality, highlighting the semantic relevance advantage of supervised models despite similar edit counts.
By comparing ranking metrics across the two dataset splits, we can easily observe that best results overall were achieved in the DENSE split. This is an expected behavior, since GNNs require interconnections for message passing procedures, i.e. to exchange and aggregate information from their neighbors.
Performance vs dataset size
Although unsupervised GAEs lag behind supervised models, there is a notable discrepancy in the required ground truth data. Supervised GNNs necessitate training on $\sim N^{2}/2$=70K pairs, maintaining a quadratic relationship with input data, entailing computationally expensive ground truth GED calculations. Conversely, unsupervised GAEs offer a more scalable solution, requiring $N$ training samples without optimizing on ground truth GED. This enables faster calculation of counterfactual explanations. While increasing training samples can enhance GAE performance, it may not be feasible in low-resource scenarios or where manual graph construction is required. Both approaches demonstrate computational efficiency, with GAE training taking less than 2 minutes for $N$=500 graphs, whereas supervised GNN training requires $\sim$3 hours for $\sim N^{2}/2$=70K graphs, with slight variations based on graph density. Despite differing training times, the retrieval, inference, and edit path computation times remain consistent between the two approaches. Training on a single GPU available in online platforms e.g. Kaggle or Google Colab is feasible for both supervised and unsupervised GNNs, while retrieval and inference operations are CPU-based, ensuring reproducibility of reported results.
4.3 Qualitative Results
Qualitative analysis is indispensable, particularly in the visual domain, where numerical metrics may not fully capture model performance. For graph retrieval, structural and semantic cues play vital roles. In Figure 3, we present results from top-performing models: supervised GCN, unsupervised ARVGA-GIN, and WL-kernel. Note that counterfactual classes do not need to align between models.
Figure 3: Counterfactuals from the best supervised GNN, unsupervised GNN and kernel for VG-DENSE (top 3) and VG-RANDOM (bottom 2).
In the case of VG-DENSE, multiple instances exhibit concordance among models, owing to its dense interconnections and the inherent ability of all models to recognize matching structures. The 1st row demonstrates unanimous agreement, as the obtained scene graph essentially forms a subgraph of the query. A human could also agree on the semantic closeness between counterfactual instances: Semantics such as ”table”, ”cup”, ”phone” are preserved, while the kitchen-related ”cake” and ”newspaper” should be edited to allow workstation-related semantics, such as ”laptop”. Moving to the 2nd row, GNN methods align; successfully integrating structural and semantic information. Despite the kernel-retrieved result having a more similar structure, GNN methods successfully discern and retrieve an image featuring a cat on furniture, surpassing a mere depiction of a bed. In terms of human perception, counterfactuals depicting a ”bedroom” (query) and a ”living room” instance (GNN-based retrievals) respect the change of class label (”bedroom”$\rightarrow$”living room”) with minimally altering semantics (”pillow” and ”cat” semantics are preserved), allowing effortless high-level interpretability. Conversely, in the 3rd row with a widespread disagreement, the merits of the supervised GNN network become evident. The unsupervised GAE captures an image resembling the query with a teddy bear wearing a bow but overlooks the presence of a dog. In contrast, the supervised GNN, having more accurately approximated GED, retrieves an image centered on a dog and a different type of toy. Interestingly, the kernel fails and retrieves an image of a train, emphasizing that a sole focus on structure (star subgraph) is insufficient.
Assessing retrieved VG-RANDOM results proves challenging due to the dissimilar nature of the sparser underlying graphs. In the 4th row, the supervised GCN retrieves a structurally closer scene graph, despite visual similarities in GNN-based counterfactuals. It becomes harder even for humans to evaluate their proximity, since all instances contain ”chair” and ”curtain” concepts, but no other discriminative structure or detail. In the final row the models reach consensus, successfully retrieving an image with palm trees, a characteristic the kernel fails to achieve despite the retrieval of very closely aligned graphs. The effectiveness of the framework relies on rich interconnections observed in VG-DENSE, underscoring the significance of expressive scene graphs, human intuition, and thoughtful data preprocessing for enhancing explanation informativeness.
Notably, none of the retrieved instances are mislead by purely visual characteristics, such as color, contrast or brightness of images. This is because graph-based algorithms naturally disregard any such characteristics, as they are not integrated into scene graphs. Using the kitchen scenario as an example, the preservation of its identity through alterations like converting to grayscale or adjusting brightness suggests that graph machine learning models, particularly GNNs, can closely align with human perception.
5 Conclusion
In this work, we unify graph-based conceptual counterfactual explanation algorithms under a common framework, upon which we conduct a comparative study. Various graph-based algorithms, such as kernels, supervised and unsupervised GNNs, highlight the value of graph representations for counterfactuals, offering insights on what needs to be changed conceptually to transit to another class efficiently. Quantitative and qualitative analysis showcases the different criteria that influence the retrieval of counterfactual instances from varying graph algorithms, and the importance of well-defined and informative semantics towards human-interpretable explanations.